OpenStudy (anonymous):

What is the solution of the equation? 4(y + 2) = 32 A. 4 B. 6 C. –10 D. 10

OpenStudy (luigi0210):

distribute the 4 out

OpenStudy (anonymous):

4y+8=32 can you solve from there ?

OpenStudy (anonymous):

Answer Is B

OpenStudy (anonymous):

4y+8=32 -8 -8 -------- 4y=24 y=6